A project that provides annual cash flows of $12,300 for 7 years costs $59,882 today.

a. If the required return is 7 percent, what is the NPV for this project?

b. Determine the IRR for this project.
Answer #1

a.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=12,300[1-(1.07)^-7]/0.07

=12,300*5.3892894

=66288.26

NPV=Present value of inflows-Present value of outflows

=66288.26-59,882

=$6406.26(Approx).

b.

Let irr be x%
At irr,present value of inflows=present value of outflows

59,882=12,300/1.0x+12,300/1.0x^2+..................+12,300/1.0x^7

Hence x=irr=10%